Flexible Budgeting and Variance Analysis I Love My Chocolate Company makes dark chocolate and light chocolate. Both productsCocoa 140,500 $4.60 0.55 Sugar 186,400 Actual Labor Hours Used Dark chocolate Actual Labor Rate $14.20 per hr. 14.80 per hr.
Answer #1
Cocoa
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($4.50/ib. - $4.60/ib ) × 140500 ib. = $14050 U F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((10 ibs. X 5300 Unit)+(7ibsX12400)-140500)X $4.50= $3150 U F
For Sugar
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($0.60 - $0.55 ) × 186400 = $9320 F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((8 ibs. X 5300 Unit)+(12 ibsX12400)-186400)X $0.60= $2880 F
Total Material Price Variance= $14050 UF+$9320 F= $4730 UF
Total Material Quantity Variance= $3150 UF+ $2880 F=$270 UF
Total Material Cost Variance= $4730 UF+ $270 UF= $500UF
Dark Choclate
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($14.5/hour - $14.20/hour) × 1930 Hours= $579 F
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard overhead recovery rate
= (0.40 Hour × 5300 Unit - 1930) × $14.50 = $2755 F
Light Choclate
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($14.50/hour - $14.80/hour) × 6350 Hours= $1905 UF
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard overhead recovery rate
= (0.50 Hour × 12400 Unit - 6350) × $14.5 = $2175 U F
Total LAbour Rate Variance= $579 F+ $1905 UF= $1326 UF
Total Labour Efficiency Variance= $2755F+$2175 UF=$580 F
Total Labour Cost Variance = $1326 UF+ $580F=$746 UF